At the end of 2017, Ayayai Corp. has accounts receivable of $1.4 million and an allowance for doubtful accounts of $85,000. On January 16, 2018, Ayayai determined that its $17,000 receivable from Blossom Company will not be collected, and management has authorized its write off. On January 31, 2018, Ayayai received notification that the company will be receiving $0.15 for every $1.00 of accounts receivable relating to Sunland Company The company had previously written off 100% of the amount due from Sunland ($55,000).

Prepare the journal entry for Ayayai Corp. to write off the Blossom receivable and any journal entry necessary to reflect the notice regarding Sunland Company

What is the estimated net realizable value of Ayayai’s accounts receivable before and after the entries above? What is the book value of Ayayai’s accounts receivable before and after the entries above?
Answer #1
Journal entries:
Date Account titles and explanation Debit Credit
2018
Jan 16. Allowance for doubtful accounts 17000
Accounts receivable 17000
(Accounts receivable written-off)
Jan 31. Accounts receivable (55000*0.15) 8250
Allowance for doubtful accounts 8250
(Accounts receivable written-off reversed)
Book value of accounts receivables:
Before entry
Book value of accounts receivables=$1400000
After entry
Book value of accounts receivables=1400000-17000+8250=$ 1391250
Net realizable value of accounts receivable=Book value of accounts receivable-Book value of allowance for doubtful accounts
Before entry
Net realizable value of accounts receivable=1400000-85000=$ 1315000
After entry
Book value of accounts receivables=$ 1391250
Book value of Net realizable value=85000-17000+8250=$ 76250
Net realizable value of accounts receivable=1391250-76250=$ 1315000